Josef's father is seven times as old as he is. One year ago, he was nine times as old as Josef. Find their age
Answer by ikleyn(52788)

.

Let x be the Joseph's age.

Then the father's age is 7x years.



One age ago 


    7x-1 = 9*(x-1)


Simplify and solve


    7x - 1 = 9x - 9

    9 - 1 = 9x - 7x

    8 = 2x.

    x = 8/2 = 4.


ANSWER.  4 years and 28 years.